the perimeter of a triangle is 32 centimeters. The first side is 3cm longer than the second side, and the third side is 6 cm shorter than three times the second side.
Answer by Paul(988)

LEt the second side be x
Let the first side be x+3
Let the third side be 3x-6
Perimeter = 32
(x)+(x+3)+(3x-6)=32
5x-3=32
5x=35
x=7 (second side)
7+3=10 (first side)
3(7)-6=15 (third side)

Hence, the three sides, are 7cm, 10cm, and 15cm.
PAUL.